# Minor DPO reject penalty to increase training robustness

### v1

@@ 3.3 DPO shortage
$L_{DPOP}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}[{log\sigma(\beta log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})}-\lambda max(0,log\frac{\pi_{ref}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)})]$ (5)
It’s same as adding an additional SFT loss on $y_{w}$, but somehow use a non-linear transformation with $\lambda max(0,log\frac{\pi_{ref}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)})$ format. It adds SFT loss only when $\pi_{\theta}(y_{w}|x)$ is smaller than $\pi_{ref}(y_{w}|x)$ so that $log\frac{\pi_{ref}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)}\geq 0$ It works like a compensation for the over penalty to the reject samples. The additional loss shall bring $rewards/chosen$ $log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}\geq 0$ for some data distribution. while it also brings an additional hyper parameter that needs to be tuned with the corresponding $\beta$ value.
@@ 3.4 Minor DPO reject penalty to improve training robustness
    For those reject samples, $\pi_{\theta}$ is more close to $\pi_{ref}$ as we stop the gradient descent on rejects samples that are $\pi_{\theta}(y_{l}|x)\leq\pi_{ref}(y_{w}|x)$, as we don’t want to over-optimized on the rejected samples
@@
Comparing to DPO, MinorDPO decreases the reject penalty and slightly increases chosen reward for samples that are $\pi_{\theta}(y_{l}|x)\leq\pi_{ref}(y_{w}|x)$ It ease the over-penalty problem and doesn’t introduce new hyper-parameter. In most cases, it can use the same DPO hyper-parameter settings(or a bigger learning rate for training). We find that minor DPO is much more robust and can accept a much higher learning rate without crashing the model. (A note here is that with higher learning rate, bigger batch size is needed to maintain the training stability)

### v3

@@ 3.3 DPO shortage
$L_{DPOP}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}[log\sigma(\beta log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}-\beta\lambda max(0,log\frac{\pi_{ref}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)}))]$ (5)
It’s same as adding an additional SFT loss on $y_{w}$, but somehow use a non-linear transformation with $\lambda max(0,log\frac{\pi_{ref}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)})$ format. It adds SFT loss only when $\pi_{\theta}(y_{w}|x)$ is smaller than $\pi_{ref}(y_{w}|x)$ so that $log\frac{\pi_{ref}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)}\geq 0$ It works like a compensation for the over penalty to the reject samples. The additional loss shall bring $rewards/chosen$ $log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}\geq 0$ for some data distribution. while it also brings an additional hyper parameter $\lambda$ that needs to be tuned with the corresponding $\beta$ value.
@@ 3.4 Minor DPO reject penalty to improve training robustness
    For those reject samples, $\pi_{\theta}$ is more close to $\pi_{ref}$ as we stop the gradient descent on rejects samples that are $\pi_{\theta}(y_{l}|x)\leq\pi_{ref}(y_{l}|x)$, as we don’t want to over-optimized on the rejected samples
@@
Comparing to DPO, MinorDPO decreases the reject penalty and slightly increases chosen reward for samples that are $\pi_{\theta}(y_{l}|x)\leq\pi_{ref}(y_{l}|x)$ It ease the over-penalty problem and doesn’t introduce new hyper-parameter. In most cases, it can use the same DPO hyper-parameter settings(or a bigger learning rate for training). We find that minor DPO is much more robust and can accept a much higher learning rate without crashing the model. (A note here is that with higher learning rate, bigger batch size is needed to maintain the training stability)